You are here: Foswiki>Commissioning Web>Logs>22Jan013_Log (23 Jan 2013, JuanGuerra)Edit Attach

UT 20130121 AO#1

Observer:Esposito,S(Arcetri),Bussoni,L(Arcetri),Puglisi,A(Arcetri),Pinna,E(Arcetri),JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO)
Telescope Operator: Steve..(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO),JChristou(LBTO),DMiller(LBTO))
Telescope Support: Kevin N(LBTO)
Instrument Support: Tim S (LBTO)

Summary

RetroRefrlector on AO#1 (AGW#2). Checking the AO Operating technical document. Performance check. Understanding how to correct the re-rotator angles respect to the new rotator position. Test of bin 1-4, the bin 4 probably needs a digital shitf of the pupils. Reset of the Wavefront sensor system system . issue: found a second spot in the psf. rotator went failure.22:30 and reboot of the jet. reset of the system.

Hand over 5:30pm

Detail

The RR is installed earluy in teh morning. The Lucifer team had access to the dome to work on nthe helium pipe, they release the telescope at 9am.

The spot shows the cross taht it tell us that the spot is wrong, because shows

Performance data

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20130122_200015 ??? 7.5 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 620 ON 56.1 H 20 500 OFF zero.fits  
20130122_200106 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 622 ON 50.5 H 20 500 OFF zero.fits  
20130122_200152 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 621 ON 60.6 H 20 500 OFF zero.fits  
20130122_200335 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 633 ON 62.3 H 20 500 OFF zero.fits  
20130122_200422 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 628 ON 62.4 H 20 500 OFF zero.fits  
20130122_200510 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.1 0.2 -1.0 3 624 ON 64.1 H 20 500 OFF zero.fits  
20130122_202100 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 622 ON 67.2 H 20 500 OFF zero.fits  
20130122_202148 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 623 ON 66.3 H 20 500 OFF zero.fits  
20130122_202236 ??? 7.5 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 620 ON 69.6 H 20 500 OFF zero.fits  
20130122_202420 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 635 ON 69.7 H 20 500 OFF zero.fits  
20130122_202505 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 624 ON 65.9 H 20 500 OFF zero.fits  
20130122_202554 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 624 ON 69.8 H 20 500 OFF zero.fits - long stop of 305 fraand impossible to continuemes at frame 1494
20130122_203050 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 627 ON 68.5 H 20 500 OFF zero.fits  
20130122_203137 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 626 ON 66.1 H 20 500 OFF zero.fits  
20130122_203224 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 630 ON 53.4 H 20 500 OFF zero.fits  
20130122_203409 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 638 ON 68.2 H 20 500 OFF zero.fits - Pupils not on target
20130122_203457 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 630 ON 66.3 H 20 500 OFF zero.fits  
20130122_203545 ??? 7.4 90 -2147483648 NaN -1.00 092604 1 400 990 0.3 0.3 0.3 3 628 ON 68.1 H 20 500 OFF zero.fits  

Screenshotpsf_secondspot.png

Screenshot_psf_offset.png

changing filter to saee if it something with the filter... Screenshot_psf_h_J.png

22:30 the rotator went in failure and the AO procedures are stopped.

23:50 With the system at close loop the wavefront sensor went in power off and with a continous failure to power on the system, There isn't a clue of who is causing this failure, but four potential candidates are; ts8, power supply, cabling of the power supply or the ethernet.

After power cycle off/on the unit removing the main power supply, the W#Unit come up, but we can an intermitent power off/on intermitent.

-- JuanGuerra - 22 Jan 2013
  • psf H and J:
    Screenshot_psf_h_J.png

  • wrong spot:
    Screenshot_psf_wrongspot.png
Topic revision: r2 - 23 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback